# BSAFusion: A Bidirectional Stepwise Feature Alignment Network for Unaligned Medical Image Fusion

### v1

@@ Abstract
If unaligned multimodal medical images can be simultaneously aligned and fused using a single-stage approach within a unified processing framework, it will not only achieve mutual promotion of dual tasks but also help reduce the complexity of the model. However, the design of this model faces the challenge of incompatible requirements for feature fusion and alignment; specifically, feature alignment requires consistency among corresponding features, whereas feature fusion requires the features to be complementary to each other. To address this challenge, this paper proposes an unaligned medical image fusion method called Bidirectional Stepwise Feature Alignment and Fusion (BSFA-F) strategy. To reduce the negative impact of modality differences on cross-modal feature matching, we incorporate the Modal Discrepancy-Free Feature Representation (MDF-FR) method into BSFA-F. MDF-FR utilizes a Modality Feature Representation Head (MFRH) to integrate the global information of the input image. By injecting the information contained in MFRH of the current image into other modality images, it effectively reduces the impact of modality differences on feature alignment while preserving the complementary information carried by different images. In terms of feature alignment, BSFA-F employs a bidirectional stepwise alignment deformation field prediction strategy based on the path independence of vector displacement between two points. This strategy solves the problem of large spans and inaccurate deformation field prediction in single-step alignment. Finally, Multi-Modal Feature Fusion block achieves the fusion of aligned features. The experimental results across multiple datasets demonstrate the effectiveness of our method. The source code is available at https://github.com/slrl123/BSAFusion/.
@@ Introduction
Multimodal medical image fusion (MMIF) involves the integration of medical image data from different imaging modalities (such as CT, MRI, PET, etc.) to create a new image that contains more comprehensive and accurate lesion information. This technology is of great significance in improving diagnostic accuracy, assisting in the development of treatment plans, promoting medical research and education, and optimizing the utilization of medical resources. As a result, it has garnered the attention of researchers, and a multitude of effective fusion algorithms have been proposed (Li et al. 2018; Tang et al. 2022b; Mu et al. 2023). However, most current methods assume that the source images being fused are strictly aligned at the pixel level. The fusion algorithm can produce the expected results only when this assumption holds true. In real scenarios, however, this assumption is often invalid. To address this issue, registration algorithms are typically used to first align the images to be fused, followed by the fusion process (as shown in Fig.1(a)). Although this two-stage method is effective, cross-modal image registration still faces numerous challenges due to differences in modalities and inconsistencies in features between images.
@@
In recent years, researchers have begun exploring the integration of multi-source image registration and fusion into a unified framework to address the aforementioned issues. By leveraging the supervision of fusion results, registration performance can be improved. Based on this idea, joint processing frameworks for image registration and fusion, such as ReCoNet (Huang et al. 2022b), UMF-CMGR (Wang et al. 2022), SuperFusion (Tang et al. 2022a), and RFNet(Xu et al. 2022), have emerged in recent years. However, these methods are not specifically designed for the registration and fusion of multimodal medical images. Although MURF (Xu, Yuan, and Ma 2023) attempts to integrate multiple types of source image fusion problems into one framework, this multitasking approach often sacrifices performance in single-task image fusion. In view of this, PAMRFuse (Zhong et al. 2023) adopts a similar approach to UMF-CMGR, focusing on the fusion of unregistered multimodal medical images. However, this method relies on images of one modality to generate corresponding images in other modalities, addressing the issue of modality differences during the registration process. Although this approach has shown some effectiveness, its registration performance is often limited by the quality of the generated images.
@@
In addition, the above methods often adopt a two-stage processing mode (as shown in Fig.1(b)), with registration preceding fusion. This approach typically necessitates the use of a separate, fully-developed image registration model. This is because registration and fusion have incompatible feature requirements, making it challenging to seamlessly embed both into the fusion process through a shared feature encoder. To tackle this issue, single-stage unaligned fusion methods, such as IVFWSR (Li et al. 2024) and RFVIF (Li et al. 2023a), have been proposed for infrared-visible images. However, these methods only address feature misalignment caused by rigid transformations and are ineffective in handling elastic transformations. They also show limitations when applied to the fusion of multimodal medical images. In fact, achieving registration and fusion of multimodal medical images using a single-stage processing mode within a joint framework remains challenging. These challenges mainly involve resolving the conflicting requirements of feature extraction for registration and fusion, and effectively eliminating the impact of modality differences on registration while preserving complementary information. Typically, feature fusion expects the features to be complementary, while feature matching demands consistency between corresponding features. To achieve simultaneous feature alignment and fusion within a single-stage processing mode, it is crucial to address the aforementioned issues.
@@
Therefore, this paper proposes a single-stage framework for multimodal medical image registration and fusion (as shown in Fig. 1(c)). Unlike traditional two-stage methods, this approach does not require a separate and complete registration process. Instead, it seamlessly embeds the registration steps into the image fusion process, effectively mitigating the increase in model complexity that would result from introducing multiple independent feature encoders. Technically, we innovatively develop an unaligned medical image fusion method called Bidirectional Stepwise Feature Alignment (BSFA). To effectively mitigate the adverse effects of modality differences on cross-modal feature matching, we integrate the Modality Discrepancy-Free Feature Representation (MDF-FR) method into the BSFA framework. MDF-FR achieves global feature integration by appending a Modality Feature Representation Head (MFRH) to each input image. This method significantly reduces the impact of modality differences and inconsistent multimodal information on feature alignment by injecting the head information of the current image into the features of the other images to be fused. As a result, this design effectively preserves the complementary information carried by different images, ensuring both the integrity and diversity of the data. For feature alignment, we propose a bidirectional stepwise deformation field prediction strategy based on the path independence of vector displacement between two points. This strategy effectively addresses the challenges of large-span and inaccurate deformation field predictions encountered in traditional single-step alignment methods, significantly enhancing both the accuracy and efficiency of feature alignment. Finally, through the Multi-Modal Feature Fusion (MMFF) module, the predicted deformation field is applied to multimodal features, achieving precise alignment and effective fusion of input images at the feature level.
@@
  • We propose a modality discrepancy reduction method. This method achieves global feature integration by appending a Modal Feature Representation Head to each input image. By incorporating the feature representation of the current image into the features of the other images to be fused, it effectively mitigates the impact of modality differences and inconsistent multimodal image information on feature alignment, while preserving complementary information across different images.
  • Based on the path independence of vector displacement between two points, a bidirectional stepwise deformation field prediction strategy is proposed. This strategy effectively addresses the challenges of large spans and inaccurate deformation field predictions encountered in traditional single-step alignment methods, significantly enhancing the accuracy and efficiency of feature alignment.
@@
For MMIF, deep learning has been widely used due to its ability to effectively extract statistical information from large datasets. Based on the types of feature extraction networks used, existing medical image fusion methods can be classified into CNN-based, Transformer-based, and hybrid methods. Among them, CNN-based methods mainly focus on network architecture design. Commonly used frameworks for MMIF in CNN-based methods include residual connections (Gu et al. 2024), skip connections (Di et al. 2024), dense connections (Zuo, Zhang, and Yang 2021), and Network Architecture Search (Ye et al. 2023). Additionally, there are dynamic meta-learning method (Huang et al. 2022a) and medical semantic-guided two-branch method (Wen et al. 2023). However, these methods are often limited by the shortcomings of CNNs in modeling long-distance dependencies. Transformer (Vaswani et al. 2017) has addressed this limitation, resulting in methods such as FATMusic (Zhao et al. 2023b) and MATR (Tang et al. 2022b). Given the complementary strengths of CNN and Transformer in feature extraction, researchers have proposed hybrid methods like DesTrans (Song et al. 2024), DFENet (Li et al. 2023b) and MRSC-Fusion (Xie et al. 2023).
In recent years, various methods have been developed for multimodel image fusion, including MMIF, such as U2Fusion (Xu et al. 2020), Cddfuse (Zhao et al. 2023a), DDFM (Zhao et al. 2023b), EMMA (Zhao et al. 2024), QuadzBayer(Zheng et al. 2024) and HFT(Chen et al. 2023). Although these methods are effective, they all assume that the source images to be fused have already been registered. However, in practical applications, this assumption often does not hold true. Therefore, when dealing with unaligned multimodal medical images, these methods cannot be directly applied and require additional image registration models to align the images for fusion. This not only increases model complexity, hindering deployment in computationally constrained environments, but also results in fusion failures if the registration model fails. To address these issues, methods for joint registration and fusion have been developed. Typical examples include ReCoNet (Huang et al. 2022b), UMF-CMGR (Wang et al. 2022), SuperFusion (Tang et al. 2022a), as well as others like RFNet (Xu et al. 2022), MURF (Xu, Yuan, and Ma 2023), IVFWSR (Li et al. 2024), and MERF (Hong, Zhang, and Ma 2024). However, these methods are not specifically designed for multimodal medical images and do not exhibit the expected advantages in this domain. Although PAMRFuse (Zhong et al. 2023) is designed for multimodal medical image fusion, its performance is limited by the quality of the generated images. To overcome these challenges, we propose an unaligned multimodal medical image fusion scheme that integrates registration and fusion, allowing the two tasks to complement each other within a single-stage framework.
@@
Dataset and Implementation Details: We follow the protocols of existing methods (Huang et al. 2024; Wen et al. 2023) and train the model using CT-MRI, PET-MRI, and SPECT-MRI datasets from Harvard [^1] http://www.med.harvard.edu/aanlib/. These datasets consist of 144, 194, and 261 strictly registered image pairs, respectively, each with a size of 256$\times$256. To simulate misaligned image pairs as collected in real-world scenarios, we designate the MRI images as the reference and apply a mixture of rigid and non-rigid deformations to the non-MRI images, thereby creating the required training set. Additionally, the same deformations are applied to 20, 55, and 77 strictly registered image pairs to construct an unaligned test set. To augment the data, we apply these mixed deformations randomly in each epoch, along with random rotations and flips to increase the diversity of the training samples.
@@
Evaluation Metrics: To objectively evaluate the performance of fusion methods, we selected five commonly used image quality metrics: Gradient-based Fusion Performance ($Q_{AB/F}$) (Xydeas, Petrovic et al. 2000), Chen-Varshney Metric ($Q_{CV}$) (Chen and Varshney 2007), Visual Information Fidelity ($Q_{VIF}$) (Han et al. 2013), Structure-based Metric ($Q_{S}$) (Piella and Heijmans 2003), and Structural Similarity Index Measure ($Q_{SSIM}$) (Wang et al. 2004). Compared to $Q_{S}$, $Q_{SSIM}$ can additionally measure brightness and contrast information. Among these evaluation metrics, a lower value of $Q_{CV}$ indicates better quality of the fused image, while higher values of the other metrics indicate better fusion quality.
Figure 5: Comparison of objective evaluation results: Joint registration and fusion method vs. the proposed method. The black line denotes the median, and the red line denotes the mean.
@@
The common approach to solving the problem of unaligned multi-source image fusion is to first perform registration on the images to be fused, obtaining aligned images, and then fuse them. We refer to this method as “Registration+Fusion.” In addition to conventional methods, a two-stage approach has been developed that combines registration and fusion into a single process, referred to as “Joint Registration and Fusion.” To verify the superiority of our proposed method, we compare it with these two approaches. Since our method is most closely related to “Joint Registration and Fusion,” we focus on the comparison with this method here. Due to space limitations, the comparison between our method and the “Registration+Fusion” method is included in the supplementary materials.
Specifically, we compared the performance of our method with five joint registration and fusion methods: UMF-CMGR, SuperFusion, MURF, IMF (Wang et al. 2024), and PAMRFuse. The first four methods are specifically designed for the registration and fusion of multimodal images and are applicable to multimodal medical image fusion. PAMRFuse, on the other hand, is a method specifically proposed for the fusion of unregistered medical images and remains the only method in this field to date. Fig. 4 presents a visual comparison of the fusion results generated by different methods. It is evident that our proposed method demonstrates significant advantages in feature alignment, contrast preservation, and detail retention. This indicates that, compared to existing two-stage joint processing frameworks, our method exhibits stronger performance, primarily due to its ability to seamlessly integrate registration and fusion tasks into a unified process. Additionally, we created box plots of test metrics for each method to visually analyze performance differences. As shown in Fig. 5, our method achieved the best mean performance across all metrics. Compared to the IMF method, which also demonstrates excellent performance, our bidirectional alignment strategy yielded significantly better fusion results, outperforming IMF’s unidirectional alignment strategy. This clearly demonstrates the effectiveness and superiority of our proposed method.
@@
Effectiveness of MDF-FR: To verify the effectiveness of MDF-FR, we designed two experimental settings: Setting A and Setting B. Both settings were aimed at demonstrating the impact of MDF-FR on our method’s performance. In Setting A, neither MFRH swapping nor the ${{\cal L}_{ce2}}$ was used.In contrast, Setting B did not involve MFRH swapping but did use the ${{\cal L}_{ce2}}$. The experimental results, as shown in Fig. 6(a), indicate that the proposed method achieves better alignment and fusion performance when MDF-FR is included.
Effectiveness of BSFA: To verify the effectiveness of BSFA, we conducted several experiments. First, we removed BSFA entirely to assess its impact on overall alignment performance and fusion results. Next, we removed the FRL from BSFA, retaining only the RRL. Finally, we kept only the FRL and removed the RRL. The alignment results shown in Fig. 6(b) indicate that the proposed method achieves excellent alignment only when BSFA is fully implemented. Please note that more detailed ablation experiments and the effectiveness of Stepwise Alignment and FusionBLK are discussed in the supplementary materials.

### v2

@@ Abstract
If unaligned multimodal medical images can be simultaneously aligned and fused using a single-stage approach within a unified processing framework, it will not only achieve mutual promotion of dual tasks but also help reduce the complexity of the model. However, the design of this model faces the challenge of incompatible requirements for feature fusion and alignment. To address this challenge, this paper proposes an unaligned medical image fusion method called Bidirectional Stepwise Feature Alignment and Fusion (BSFA-F) strategy. To reduce the negative impact of modality differences on cross-modal feature matching, we incorporate the Modal Discrepancy-Free Feature Representation (MDF-FR) method into BSFA-F. MDF-FR utilizes a Modality Feature Representation Head (MFRH) to integrate the global information of the input image. By injecting the information contained in MFRH of the current image into other modality images, it effectively reduces the impact of modality differences on feature alignment while preserving the complementary information carried by different images. In terms of feature alignment, BSFA-F employs a bidirectional stepwise alignment deformation field prediction strategy based on the path independence of vector displacement between two points. This strategy solves the problem of large spans and inaccurate deformation field prediction in single-step alignment. Finally, Multi-Modal Feature Fusion block achieves the fusion of aligned features. The experimental results across multiple datasets demonstrate the effectiveness of our method. The source code is available at https://github.com/slrl123/BSAFusion/.
@@ Introduction
Multimodal medical image fusion (MMIF) involves the integration of medical image data from different imaging modalities (such as CT, MRI, PET, etc.) to create a new image that contains more comprehensive and accurate lesion information. This technology is of great significance in improving diagnostic accuracy, assisting in the development of treatment plans, promoting medical research and education, and optimizing the utilization of medical resources. As a result, it has garnered the attention of researchers, and a multitude of effective fusion algorithms have been proposed (Li et al. 2018; Tang et al. 2022b; Mu et al. 2023a). However, most current methods assume that the source images being fused are strictly aligned at the pixel level. The fusion algorithm can produce the expected results only when this assumption holds true. In real scenarios, however, this assumption is often invalid. To address this issue, registration algorithms are typically used to first align the images to be fused, followed by the fusion process (as shown in Fig.1(a)). Although this two-stage method is effective, cross-modal image registration still faces numerous challenges due to differences in modalities and inconsistencies in features between images.
@@
In recent years, researchers have begun exploring the integration of multi-source image registration and fusion into a unified framework to address the aforementioned issues. By leveraging the supervision of fusion results, registration performance can be improved. Based on this idea, joint processing frameworks for image registration and fusion have emerged in recent years (Huang et al. 2022b; Wang et al. 2022; Tang et al. 2022a; Xu et al. 2022). However, these methods are not specifically designed for the registration and fusion of multimodal medical images. Although MURF (Xu, Yuan, and Ma 2023) attempts to integrate multiple types of source image fusion problems into one framework, this multitasking approach often sacrifices performance in single-task image fusion. In view of this, PAMRFuse (Zhong et al. 2023) adopts a similar approach to UMF-CMGR, focusing on the fusion of unregistered multimodal medical images. However, this method depends on an image of one modality to generate a corresponding image in another modality, and its registration performance is often constrained by the quality of the generated image.
@@
In addition, the above methods often adopt a two-stage processing mode (as shown in Fig.1(b)), with registration preceding fusion. This approach typically necessitates the use of a separate, fully-developed image registration model. This is because registration and fusion have incompatible feature requirements, making it challenging to seamlessly embed both into the fusion process through a shared feature encoder. To tackle this issue, single-stage unaligned fusion methods, such as IVFWSR (Li et al. 2024a) and RFVIF (Li et al. 2023a), have been proposed for infrared-visible image fusion. However, these methods only address feature misalignment caused by rigid transformations and are ineffective in handling elastic transformations. In fact, achieving registration and fusion of multimodal medical images using a single-stage processing mode within a joint framework remains challenging. These challenges mainly involve resolving the conflicting requirements of feature extraction for registration and fusion. Typically, feature fusion expects the features to be complementary, while feature matching demands consistency between corresponding features. To achieve simultaneous feature alignment and fusion within a single-stage processing mode, it is crucial to address the aforementioned issues.
@@
Therefore, this paper proposes a single-stage framework for multimodal medical image registration and fusion (as shown in Fig. 1(c)). Unlike traditional two-stage methods, this approach does not require a separate and complete registration process. Instead, it seamlessly embeds the registration steps into the image fusion process, effectively mitigating the increase in model complexity that would result from introducing multiple independent feature encoders. Technically, we innovatively develop an unaligned medical image fusion method called Bidirectional Stepwise Feature Alignment (BSFA). To effectively mitigate the adverse effects of modality differences on cross-modal feature matching, we integrate the Modality Discrepancy-Free Feature Representation (MDF-FR) method into the BSFA framework. MDF-FR achieves global feature integration by appending a Modality Feature Representation Head (MFRH) to each input image. This method significantly reduces the impact of modality differences and inconsistent multimodal information on feature alignment by injecting the head information of the current image into the features of the other images to be fused. As a result, this design effectively preserves the complementary information carried by different images, ensuring both the integrity and diversity of the data. For feature alignment, we propose a bidirectional stepwise deformation field prediction strategy based on the path independence of vector displacement between two points. This strategy effectively addresses the challenges of large-span and inaccurate deformation field predictions encountered in traditional single-step alignment methods, significantly enhancing both the accuracy and efficiency of feature alignment. Finally, through the Multi-Modal Feature Fusion (MMFF) module, the predicted deformation field is applied to multimodal features, achieving precise alignment and effective fusion of input images at the feature level. Overall, the contributions of this paper can be summarized as follows:
@@
  • We propose a modality discrepancy reduction method. This method achieves global feature integration by appending an MFRH to each input image. By incorporating the feature representation of the current image into the features of the other images to be fused, it effectively mitigates the impact of modality differences on feature alignment.
  • Based on the path independence of vector displacement between two points, a bidirectional stepwise deformation field prediction strategy is proposed. It effectively addresses the challenges of large spans and inaccurate deformation field predictions encountered in traditional single-step alignment methods.
@@
For MMIF, deep learning has been widely used due to its ability to effectively extract statistical information from large datasets. Based on the types of feature extraction networks used, existing medical image fusion methods can be classified into CNN-based, Transformer-based, and hybrid methods. Among them, CNN-based methods mainly focus on network architecture design. Commonly used frameworks for MMIF in CNN-based methods include residual connections (Gu et al. 2024), skip connections (Di et al. 2024), dense connections (Zuo, Zhang, and Yang 2021), and Network Architecture Search (Ye et al. 2023). Additionally, there are dynamic meta-learning method (Huang et al. 2022a) and medical semantic-guided two-branch method (Wen et al. 2023a). However, these methods are often limited by the shortcomings of CNNs in modeling long-distance dependencies. Transformer (Vaswani et al. 2017) has addressed this limitation, resulting in methods such as FATMusic (Zhao et al. 2023b) and MATR (Tang et al. 2022b). Given the complementary strengths of CNN and Transformer in feature extraction (Li et al. 2024b), researchers have proposed hybrid methods like DesTrans (Song et al. 2024), DFENet (Li et al. 2023b) and MRSC-Fusion (Xie et al. 2023).
In recent years, various methods have been developed for multimodel image fusion, including MMIF, such as U2Fusion (Xu et al. 2020), Cddfuse (Zhao et al. 2023a), DDFM (Zhao et al. 2023b), EMMA (Zhao et al. 2024a), QuadzBayer(Zheng et al. 2024) and HFT(Chen et al. 2023). Although these methods are effective, they all assume that the source images to be fused have already been registered. However, in practical applications, this assumption often does not hold true. Therefore, when dealing with unaligned multimodal medical images, these methods cannot be directly applied and require additional image registration models to align the images for fusion. This not only increases model complexity, hindering deployment in computationally constrained environments, but also results in fusion failures if the registration model fails. To address these issues, methods for joint registration and fusion have been developed. Typical examples include ReCoNet (Huang et al. 2022b), UMF-CMGR (Wang et al. 2022), SuperFusion (Tang et al. 2022a), as well as others like RFNet (Xu et al. 2022), MURF (Xu, Yuan, and Ma 2023), IVFWSR (Li et al. 2024a), and MERF (Hong, Zhang, and Ma 2024). However, these methods are not specifically designed for multimodal medical images and do not exhibit the expected advantages in this domain. Although PAMRFuse (Zhong et al. 2023) is designed for MMIF, its performance is limited by the quality of the generated images. To overcome these challenges, we propose an unaligned MMIF scheme that integrates registration and fusion, allowing the two tasks to complement each other within a single-stage framework.
@@
Dataset and Implementation Details: We follow the protocols of existing methods (Huang et al. 2024a; Wen et al. 2023a) and train the model using CT-MRI, PET-MRI, and SPECT-MRI datasets from Harvard [^1] http://www.med.harvard.edu/aanlib/. These datasets consist of 144, 194, and 261 strictly registered image pairs, respectively, each with a size of 256$\times$256. To simulate misaligned image pairs as collected in real-world scenarios, we designate the MRI images as the reference and apply a mixture of rigid and non-rigid deformations to the non-MRI images, thereby creating the required training set. Additionally, the same deformations are applied to 20, 55, and 77 strictly registered image pairs to construct an unaligned test set. To augment the data, we apply these mixed deformations randomly in each epoch, along with random rotations and flips to increase the diversity of the training samples.
@@
Evaluation Metrics: To objectively evaluate the performance of fusion methods, we selected five commonly used image quality metrics: Gradient-based Fusion Performance ($Q_{AB/F}$) (Xydeas, Petrovic et al. 2000), Chen-Varshney Metric ($Q_{CV}$) (Chen and Varshney 2007), Visual Information Fidelity ($Q_{VIF}$) (Han et al. 2013), Structure-based Metric ($Q_{S}$) (Piella and Heijmans 2003), and Structural Similarity Index Measure ($Q_{SSIM}$) (Wang et al. 2004). Among these evaluation metrics, a lower value of $Q_{CV}$ indicates better quality of the fused image, while higher values of the other metrics indicate better fusion quality.
Figure 5: Comparison of objective evaluation results: Joint registration and fusion vs. the proposed method. The black line denotes the median, and the red line denotes the mean.
@@
The common approach to solving the problem of unaligned multi-source image fusion is to first perform registration on the images to be fused, and then fuse them. We refer to this method as “Registration+Fusion.” In addition to conventional methods, a two-stage approach has been developed that combines registration and fusion into a single process, referred to as “Joint Registration and Fusion.” To verify the superiority of our method, we compare it with these two approaches. Since our method is most closely related to “Joint Registration and Fusion,” we focus on the comparison with this method here. Due to space limitations, the comparison between our method and the “Registration+Fusion” method is included in the supplementary materials.
Specifically, we compared the performance of our method with five joint registration and fusion methods: UMF-CMGR, SuperFusion, MURF, IMF (Wang et al. 2024), and PAMRFuse. The first four methods are specifically designed for the registration and fusion of multimodal images and are applicable to MMIF. PAMRFuse, on the other hand, is a method specifically proposed for the fusion of unregistered medical images. Fig. 4 presents a visual comparison of the fusion results generated by different methods. It is evident that our proposed method demonstrates significant advantages in feature alignment, contrast preservation, and detail retention. This indicates that, compared to existing two-stage joint processing frameworks, our method exhibits stronger performance, primarily due to its ability to seamlessly integrate registration and fusion tasks into a unified process. Additionally, we created box plots of test metrics for each method to visually analyze performance differences. As shown in Fig. 5, our method achieved the best mean performance across all metrics. Compared to the IMF method, which also demonstrates excellent performance, our bidirectional alignment strategy yielded significantly better fusion results, outperforming IMF’s unidirectional alignment strategy.
@@
Effectiveness of MDF-FR: To verify the effectiveness of MDF-FR, we designed Setting A and Setting B. In Setting A, neither MFRH swapping nor the ${{\cal L}_{ce2}}$ was used. In contrast, Setting B did not involve MFRH swapping but did use the ${{\cal L}_{ce2}}$. The experimental results, as shown in Fig. 6(a), indicate that the proposed method achieves better alignment and fusion performance when MDF-FR is included. Effectiveness of BSFA: To verify the effectiveness of BSFA, we conducted several experiments. First, we removed BSFA entirely to assess its impact on overall alignment performance and fusion results. Next, we removed the FRL from BSFA, retaining only the RRL. Finally, we kept only the FRL and removed the RRL. The alignment results shown in Fig. 6(b) indicate that the proposed method achieves excellent alignment only when BSFA is fully implemented.